Tohban Report 2013-08-28

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 21 August 2013
End Date: 28 August 2013
Tohban: Mitsuo Oka
Tohban email:
Next Tohban: Juan Carlos Martinez Oliveros
List all reports


Solar Activity

Solar activity was low this week. There are currently 7 active regions on the disk, a few of which will rotate over the limb within 1-2 days. STEREO B shows several small active regions on the farside of the Sun.

How many GOES flares occurred?

 Flares above B, C, M, X class were      5    18     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      2     7     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      2     1     0     0

There were RHESSI flares/GOES flares 73 / 23 over the time range 20-Aug-13 27-Aug-13

Memory Management

The SSR is emptying by the end of the daily pass set.

Spacecraft Status

Cold plate 1 and 2 are currently at 117.1 K and 115.8 K. Cryocooler looks normal. Spin rate is 14.6 and needs to be increased soon.

Data Gaps

The data packets from Wallops are still not being processed and a number of large data gaps are visible in the browser plots.

Detector issues

Other notes

A circuit breaker tripped at the MOC during the weekend. The cause is unknown and an electrician will take a look at it.

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled, during entire eclipse
Night time data (rears) Enabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Increased noise in Detector 7 continues.
Personal tools